Calculate the perimeter of a rectangle whose area is 25x2 – 35x + 12. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student


Area of rectangle = 25x^2– 35x + 12

We know, area of rectangle = length × breadth

So, by factoring 25x^2– 35x + 12, the length and breadth can be obtained.

25x^2– 35x + 12 = 25x^2– 15x – 20x + 12

=> 25x^2– 35x + 12 = 5x(5x – 3) – 4(5x – 3)

=> 25x^2– 35x + 12 = (5x – 3)(5x – 4)

So, the length and breadth are (5x – 3)(5x – 4).

Now, perimeter = 2(length + breadth)

perimeter of the rectangle = 2[(5x – 3)+(5x – 4)]

= 2(5x – 3 + 5x – 4) = 2(10x – 7) = 20x – 14

Thus, the perimeter = 20x – 14


Thanks